# Towards Fast Inference: Exploring and Improving Blockwise Parallel Drafts

### v1

@@ 6.1 Consecutive Repetition
Table 3: Consecutive token repetition present in BPD drafts before and after C4-trained 2-gram rescoring of the top-16 lattice. “Consec” is the percentage of consecutive identical draft tokens out of all pairs of consecutive tokens. “Max Run” is the average maximum repeated subsequence length in tokens (upper bound of nine, the number of BPD heads). Higher values correspond to more egregious repetition in drafts.
@@
We observe that vanilla BPD drafts are prone to significant token repetition. This is due to the fact that each head’s prediction is independent of the others, and is a limitation shared with non-autoregressive generation in general (Gu et al., 2017). Table 3 shows the proportion of consecutive tokens in BPD drafts that are identical to each other, along with the average maximum length of repeated sequences in BPD drafts across all decode time steps. We compare these statistics before and after rescoring with a 2-gram LM: a trivial rescorer, but one that can encourage local consistency between consecutive draft tokens. Strings of repeated tokens are unnatural, and unlikely to be generated by a strong base language model. Rescoring the BPD top-$k$ lattice with even a simple language model eliminates a significant amount of repetition, reducing the percentage of consecutive repeated tokens from between 9.9% to 24.5% absolute.
@@ 6.2 Confidence Across Multiple Heads
Intuitively, predicting the probability distribution of the next $i^{\text{th}}$ token becomes harder as $i$ increases. To better understand this phenomenon, we measure the confidence of the predictions by the entropy of the probability distribution. In Figure 2, we plot the normalized histogram of entropy of each head on the LAMBADA task. From the normalized histogram, it is clear that the entropy increases as we move from first head to the last head, which agrees with our intuition that hardness of predictions increases as $i$ increases.
However, we observed that the entropy of heads does not increase monotonically for all tasks. Let ¯$\overline{\mathbb{H}}[i]$ be the average entropy of head $i$ on a particular corpus, and let $h_{\max}=\max\limits_{k}\{k:\forall i<k,\overline{\mathbb{H}}[i]\leq\overline{\mathbb{H}}[i+1]\}$ , be the index of the largest head such that the average entropy of each head increases monotonically to that point. We observed a strong correlation between $h_{\max}$ and block efficiency (Figure 3). Heads with lower entropy (indicating more confident predictions) intuitively contribute more to efficiency. Nonetheless, simply maximizing the number of low-entropy heads is not optimal, but rather incorporating progressively higher entropy heads, up to a certain point, can benefit decoding efficiency. A linear regression confirms this with an R-value of $\mathbf{0.77}$. This analysis suggests that BPD head entropy could be used as a proxy for block efficiency, and thus inference latency.
Figure 2: Entropy distributions across BPD heads on LAMBADA. The density plots illustrate the entropy distribution for each head in the model.
Figure 3: Correlation between block efficiency and $h_{\max}$, the head till which the average entropy in a task increases monotonically.
@@ 6.3 Oracle Top-$k$ Block Efficiency
Figure 4: An example of a top-5 BPD sausage lattice generated on a NewsRoom example. Edge weights correspond to (rescored) logits. Edges at each time step are ordered in descending weight and green, bolded edges correspond to candidates matching the greedy decode over the next nine tokens: ”… desktop computers with new Intel Corp processors that it …”. The initial node in this graph is state 0 and the final node is 9.
@@
Oracle efficiency
The concept of oracle block efficiency in BPD serves as a theoretical benchmark, illustrating the headroom available from improving the quality of the BPD draft. To compute oracle block efficiency, we consider the top-$k$ most probable tokens at each head, and form a “sausage” lattice from these. This data structure is a weighted directed graph, which succinctly represents all possible drafts (and their score under the BPD model) that could be formed from selecting one of $k$ tokens from each of the $h$ heads (Figure 4). In the automatic speech recognition and machine translation communities, it is known as a “confusion network” (Shen et al., 2006; Koehn et al., 2007).
Given the top-$k$ BPD lattice at each decoding step, we identify an oracle path that represents the path through the lattice that maximizes the length of the accepted prefix. This exercise, as shown in Figure 5, gives us insight into how much headroom exists in improving BPD drafts.
Potential headroom from oracle selection
Oracle drafting is not practical, but rather a reference point. Analyzing the gap between actual BPD performance and the oracle upper bound (Figure 6) helps us to understand the limitations of the original BPD drafts and potential areas for improvement. Additionally, exploring oracle efficiency as a function of the $k$ in the top-$k$ lattice, demonstrates how “close” the BPD draft was to producing a stronger draft.
@@
Figure 6: Oracle block efficiency over the BPD top-$k$ lattice as a function $k$. Each plot (a-f) represents a different task, demonstrating the relative improvement in block efficiency of the oracle draft with respect to the standard BPD draft as a function of the number of BPD heads used.
@@ 7 Lattice Rescoring for Improved Block Efficiency
Having explored BPD’s prediction dynamics, we propose two drafting algorithms to improve block efficiency through rescoring of the BPD lattice. This section presents techniques for rescoring the top-$k$ BPD lattice along with empirical results.
Each of these algorithms is a modification of the BPD drafted in Stage 1 in Algorithm 1. Instead of using the most likely token at each head as the prediction, we construct the top-$k$ sausage lattice of likely drafts from each head, where the set of top-$k$ tokens is denoted as $S_{i}$ for head $i$. This approach allows any token within $S_{i}$ to be chosen for position $i$, yielding a total possible combinations of:
@@
Algorithm 2 Local rescoring via neural models
0:  : Blockwise parallel decoder $\mathcal{M}_{\theta}^{h}$, top-$k$ indices selection function Top-$k(\cdot)$, rescoring model $\mathcal{M}_{\theta_{r}}$, interpolation weight $\alpha>0$.
1:  $z^{1}_{t+1},\dots,z^{h}_{t+h}\leftarrow\mathcal{M}_{\theta}^{h}(y_{t+1},\dots,y_{t+h}|\bar{x},y_{\leq t})$
2:  $S_{1},\dots,S_{h}\leftarrow\textsc{Top-}k(z^{1}_{t+1},\dots,z^{h}_{t+h})$
3:  /* Lattice Rescoring (subsection 7.1)*/
4:  for $j\leftarrow 2,\dots,h$ in parallel do
5:     $r_{t+j}\leftarrow\mathcal{M}_{\theta_{r}}(y_{t+j}|x,y_{<t+j})$
6:     $z^{j}_{t,j}[S_{j}]\leftarrow z^{j}_{t,j}[S_{j}]+\alpha\cdot r_{t+j}[S_{j}]$
7:  end for
Table 4: Block efficiency of various lattice rescoring methods over the top-$16$ BPD lattice. Parallel BPD variants consider the $16$ most probable rescored paths through the lattice. “Parallel Base BPD” indicates parallel BPD performance over the original BPD lattice. Relative percent improvement over Base is indicated in parentheses.
Task Dataset Base 4-gram neural-61M Parallel Base BPD Parallel 4-gram BPD Oracle (k=16)
@@ 7.1 Local Rescoring via Neural Models
A simple approach uses a small neural rescorer, interpolating between the logits of the rescorer LM and BPD logits with a tuned interpolation weight (line 3-6 in Algorithm 2). The locally rescored prediction is given by:
@@
where $z^{j}_{t,j}$ represents the BPD logit vector at head $j$, and $r_{t+j}$ is the vector of logits predicted by the small neural rescoring model, which is conditioned on the sequence $y_{\leq t},\cdots,\hat{y}_{t+2},\cdots,\hat{y}_{t+j-1}$. The parameter $\alpha$ is the weight placed on the rescorer’s prediction. We experiment with decoder-only transformers having 32, 61, and 94 million (M) weight parameters (Appendix C). We use greedy rescoring when generating the neural draft.
7.2 Parallel BPD via Global n-gram Rescoring
We also evaluate the quality of drafts generated by rescoring with an n-gram LM. Recall that BPD LMs can be used to compute a lattice representing $k^{h}$ possible sequences. We rescore all of these sequences with an n-gram model, select the top $p$ sequences and pass them to the verification stage. When $p=1$, we refer to this as n-gram rescoring and when $p>1$, we refer to this as the Parallel BPD.
While global rescoring typically yields better results compared to local rescoring, rescoring $k^{h}$ sequences with a neural LM and selecting the most likely sequence would take time $O(k^{h})$, which is computationally prohibitive in most cases. Hence, we take advantage of n-gram models, which are unique in that we can select the most likely sequence in time $\text{poly}(k,h)$ via dynamic programming. We use the OpenFST library (Allauzen et al., 2007) to represent each n-gram model as a weighted finite state automaton and apply finite state composition with the top-$k$ lattice followed by extraction of the $p$ most likely draft sequences. Training details for the n-gram models are given in Section C.3.
@@ 7.2 Parallel BPD via Global n-gram Rescoring
Figure 7: Block efficiency of parallel BPD as a function of the number of top $p$ sequences verified in parallel. Each plot (a-f) displays the block efficiency achieved with parallel BPD after n-gram -rescoring compared to the the same number of paths extracted from the original top-$16$ BPD lattice. The number of top $p$ sequences is varied in ${1,2,4,8,16}$ over a top-$16$ BPD lattice.
@@ 7.3 Empirical Evaluation
Block efficiency Table 4 and Figure 7 demonstrate the impact of lattice rescoring on block efficiency across various tasks. Autoregressive neural, n-gram LM, and parallel BPD rescoring all demonstrate improvements in block efficiency, although gains are task-dependent.
  • Tasks with high initial block efficiency (LAMBADA, CNN/Daily): Both rescoring methods show little to no improvement, suggesting that vanilla BPD already produces high quality drafts.
  • Tasks with low initial block efficiency (SQuAD V1, SAMSUM, XSUM, NewsRoom): Both neural and n-gram rescoring lead to block efficiency gains, particularly with neural LMs achieving the best performance in some cases. This suggests that rescoring helps refine predictions and navigate the lattice more efficiently in these scenarios.
Table 5: Wins, ties, and losses of 61M parameter neural-rescored drafts and vanilla BPD. “Repair” corresponds to instances where the rescored draft eliminates repetition and “Regress” corresponds to instances where the rescored draft introduces repetition.
@@
Repairing repetitions
In subsection 6.1, we note that vanilla BPD drafts are prone to token-level repetition and that rescoring with a simple language model reduces the incidence of this. Although rescoring reduces repetition overall in drafts, is this driving improvements in block efficiency? To answer this, we compared the drafts generated by greedy rescoring with the 61M parameter neural rescorer against vanilla BPD. Time step instances were considered wins/ties/losses based on the accepted prefix length of the rescored draft vs. vanilla BPD. Table 5 displays the win frequency across tasks along with the percentage of wins/losses that can be attributed to introducing/eliminating repetition.
Note that in the tasks where rescoring improves block efficiency the most, NewsRoom and MultiNews, a high percentage of those repaired instances are driven by fixing erroneously repeated tokens. In fact, for MultiNews, 66.23% of BPD drafts are improved through repetition repair.
We also evaluated the performance of rescoring with in-domain trained rescoring LMs, but found that they tended to perform no better than C4-trained LMs (Appendix D).
@@ 8 Conclusion
In this paper, we present a comprehensive analysis of BPD, highlighting its predictive dynamics and proposing methods to enhance its inference speed. Our study offers insights into BPD’s behavior, particularly the tendency for drafts to contain consecutive repetitions and its heads to exhibit varying confidence levels in predictions. We introduce a novel measure, oracle top-$k$ block efficiency, to explore potential improvements in block efficiency. Two algorithms are proposed: one using neural models for local rescoring and another employing n-gram models for Parallel BPD. These algorithms leverage the strengths of both BPD and small rescoring models to accelerate the decoding process, pushing the boundaries of efficient text generation with BPD. We believe that this paper lays the groundwork for future explorations in optimizing language models for efficient decoding speed without compromising the output quality.
Broader Impact
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted in the body. However, from a broader perspective, there are some potential implications related to systematic impact. These discussions are detailed further in the Appendix A.
@@ Appendix A Broader Impact
  • Inference Efficiency: We’ve achieved further acceleration in inference with minimal increases in CPU overhead and memory bandwidth utilization.
  • Scalability: The introduction of FST for selecting top $p$ sequences from top $k$ predictions demonstrates notable scalability, hinting at accelerated inference with growing model sizes. This scalability is further explored through diverse experiments involving various n-gram  and neural LM architectures.
  • Systematic impact: The advancement of BPD promises identical performance to traditional autoregressive methods, but with significantly improved efficiency in decoding times. In parallel with speculative decoding research, our approach can reshape the landscape of LLM, offering faster, more accessible AI solutions across various industries.
@@ B.1 Limitation
In this work, we focus on refining BPD drafts for greedy decoding, the decoding scheme considered in the original BPD paper (Stern et al., 2018). Extending this methodology to non-greedy sampling strategies is an area for future work. Leviathan et al. (2023) illuminates the feasibility of incorporating speculative decoding strategies to expedite non-greedy sampling processes. Their approach underscores the universality of the verification process across diverse model architectures, contingent on the availability of draft token-level predictions and probabilities — a requirement met by (rescored) BPD candidates.
@@
We show that lattice rescoring can improve the block efficiency for a 1.5 billion parameter BPD LM. It remains an open question as to whether one would observe similar improvements in block efficiency with larger LMs.
@@
A notable challenge in our framework is the memory overhead encountered during parallel verification. The simultaneous processing of multiple token predictions, particularly for larger models and draft batches when coupled with lattice rescoring, can intensify memory demands. That said, this can be mitigated by the application of tree attention during verification (Spector & Re, 2023).
@@
The implementation of drafting heads within our BPD framework, while foundational, leaves room for advanced development. This aspect is particularly crucial for refining the model’s accuracy and efficiency. Future iterations could benefit from more sophisticated training methods, enhancing these heads’ ability to navigate complex linguistic contexts and improving overall predictive performance.

### v2

@@ 6.1 Consecutive repetition
Table 3: Consecutive token repetition in block drafts before and after C4-trained 2-gram rescoring of the top-16 lattice. “Consec" is the percentage of consecutive identical draft tokens out of all pairs of consecutive tokens. “Max run" is the average maximum repeated subsequence length in tokens (upper bound of 9, the number of block draft heads). Higher values correspond to more egregious repetition in drafts.
@@
We observe that vanilla block drafts are prone to significant token repetition. This is due to the fact that each head’s prediction is independent of the others, and is a limitation shared with non-autoregressive generation in general [14]. Table 3 shows the proportion of consecutive tokens in block drafts that are identical to each other, along with the average maximum length of repeated sequences in block drafts across all decode time steps. We compare these statistics before and after rescoring with a 2-gram LM: a trivial rescorer, but one that can encourage local consistency between consecutive draft tokens. Strings of repeated tokens are unnatural, and unlikely to be generated by a strong base language model. Rescoring the top-$k$ lattice with even a simple language model eliminates a significant amount of repetition, reducing the percentage of consecutive repeated tokens from between 9.9% to 24.5%, depending on the task.
(a) Entropy distributions across block draft heads
(b) Correlation between block efficiency and $h_{\max}$
Figure 3: (a) Entropy distributions across block draft heads on LAMBADA. The density plots illustrate the entropy distribution for each head in the model. (b) Correlation between block efficiency and $h_{\max}$, the head until which the average entropy in a task increases monotonically.
@@ 6.2 Confidence across multiple heads
Intuitively, predicting the identity of the $i^{\text{th}}$ future token becomes harder as $i$ increases. To better understand this phenomenon, we measure the confidence of the predictions by the entropy of the probability distribution. In 3(a), we plot the normalized histogram of entropy of each head on the LAMBADA task. From the normalized histogram, it is clear that the entropy increases as we move from first head to the last head, which agrees with our intuition that hardness of predictions increases as $i$ increases.
However, we observed that the entropy of heads does not increase monotonically for all tasks. Let ¯$\overline{\mathbb{H}}[i]$ be the average entropy of head $i$ on a particular corpus, and let $h_{\max}=\max\limits_{k}\{k:\forall i<k,\overline{\mathbb{H}}[i]\leq\overline{\mathbb{H}}[i+1]\}$, be the index of the largest head such that the average entropy of each head increases monotonically to that point. We observed a strong correlation between $h_{\max}$ and block efficiency (3(b)). Heads with lower entropy (indicating more confident predictions) intuitively contribute more to efficiency. Nonetheless, simply maximizing the number of low-entropy heads is not optimal, but rather incorporating progressively higher entropy heads, up to a certain point, can benefit decoding efficiency. A linear regression confirms this with an R-value of $\mathbf{0.77}$. This analysis suggests that BPD head entropy could be used as a proxy for block efficiency, and thus inference latency.
@@ 6.3 Oracle top-k block efficiency
Figure 4: An example of a top-5 sausage lattice generated on a NewsRoom example. Edge weights correspond to (rescored) logits. Edges at each time step are ordered in descending weight and green, bolded edges correspond to candidates matching the greedy decode over the next nine tokens: "… desktop computers with new Intel Corp processors that it …". The initial node in this graph is state 0 and the final node is 9.
Oracle efficiency  The concept of oracle block efficiency in BPD serves as a theoretical benchmark, illustrating the headroom available from improving the quality of the block draft. To compute oracle block efficiency, we consider the top-$k$ most probable tokens at each head, and form a “sausage” lattice from these. This data structure is a weighted directed graph, which succinctly represents all possible drafts (and their score under the BPD model) that could be formed from selecting one of $k$ tokens from each of the $h$ heads (Figure 4). In the automatic speech recognition and machine translation communities, it is known as a “confusion network” [36, 23].
@@
Given the top-$k$ lattice at each decoding step, we identify an oracle path that represents the path through the lattice that maximizes the length of the accepted prefix. This exercise, as shown in Figure 5, gives us insight into how much headroom exists in improving block drafts.
@@
Figure 6: Oracle block efficiency over the top-$k$ lattice as a function $k$. Each plot (a-f) represents a different task, demonstrating the relative improvement in block efficiency of the oracle draft with respect to the standard block draft as a function of the number of block draft heads used.
Potential headroom from oracle selection  Oracle drafting is not practical, but rather a reference point. Analyzing the gap between actual BPD performance and the oracle upper bound (Figure 6) helps us to understand the limitations of the original block drafts and potential areas for improvement. Additionally, exploring oracle efficiency as a function of the $k$ in the top-$k$ lattice, demonstrates how “close” the block draft was to producing a stronger draft.
@@ 7 Lattice rescoring for improved block efficiency
Having explored BPD’s prediction dynamics, we propose two drafting algorithms to improve block efficiency through rescoring of the top-$k$ lattice. This section presents techniques for rescoring the top-$k$ lattice along with empirical results.
Each of these algorithms is a modification of the block drafted in Stage 1 in Algorithm 1. Instead of using the most likely token at each head as the prediction, we construct the top-$k$ sausage lattice of likely drafts from each head, where the set of top-$k$ tokens is denoted as $S_{i}$ for head $i$. This approach allows any token within $S_{i}$ to be chosen for position $i$, yielding a total possible combinations of:
@@
[…]
Task Dataset Baseline Local rescoring Global rescoring Oracle (k=16)
BPD neural-61M BPD 4-gram BPD 16-best 0-gram BPD 16-best 4-gram BPD
@@ 7.1 Local rescoring via neural models
10:  : Blockwise parallel LM $\mathcal{M}_{\theta}^{h}$, top-$k$ indices selection function Top-$k(\cdot)$, rescoring model $\mathcal{M}_{\theta_{r}}$, interpolation weight $\alpha>0$.
21:  $z^{1}_{t+1},\dots,z^{h}_{t+h}\leftarrow\mathcal{M}_{\theta}^{h}(y_{t+1},\dots,y_{t+h}|\bar{x},y_{\leq t})$
32:  $S_{1},\dots,S_{h}\leftarrow\textsc{Top-}k(z^{1}_{t+1},\dots,z^{h}_{t+h})$
43:  /* Lattice Rescoring (subsection 7.1)*/
4:  for $j\leftarrow 2,\dots,h$ in parallel do
5:    $r_{t+j}\leftarrow\mathcal{M}_{\theta_{r}}(y_{t+j}|x,y_{<t+j})$
6:    $z^{j}_{t,j}[S_{j}]\leftarrow z^{j}_{t,j}[S_{j}]+\alpha\cdot r_{t+j}[S_{j}]$
7:  end for
Algorithm 2 Local rescoring via neural models
A simple approach uses a small neural rescorer, interpolating between the logits of the rescorer LM and vanilla block draft logits with an interpolation weight (algorithm 2). The rescored prediction is given by:
@@
where $z^{j}_{t,j}$ represents the logit of the block draft at head $j$, and $r_{t+j}$ is the corresponding logit predicted by the small neural rescoring model, which is conditioned on the sequence $y_{\leq t},\cdots,\hat{y}_{t+2},\cdots,\hat{y}_{t+j-1}$. The parameter $\alpha$ is the weight placed on the rescorer’s prediction. We experiment with decoder-only transformers having 32, 61, and 94 million (M) weight parameters (Appendix C). We use greedy rescoring when generating the neural draft.
7.2 Global n-gram rescoring
We also evaluate the quality of drafts generated by rescoring with an n-gram LM. Recall that blockwise parallel LMs can be used to compute a lattice representing $k^{h}$ possible sequences. We rescore all of these sequences with an n-gram  model, select the top $p$ sequences and pass them to the verification stage. When $p=1$, we refer to this as n-gram rescoring and when $p>1$, we refer to this as $p$-best n-gram BPD.
While global rescoring typically yields better results compared to local rescoring, rescoring $k^{h}$ sequences with a neural LM and selecting the most likely sequence would take time $O(k^{h})$, which is computationally prohibitive in most cases. Hence, we take advantage of n-gram models, which are unique in that we can select the most likely sequence in time $\text{poly}(k,h)$ using dynamic programming. We use the OpenFST library [1] to represent each n-gram model as a weighted finite state automaton and apply finite state composition with the top-$k$ lattice followed by extraction of the $p$ most likely draft sequences. Training details for the n-gram models are given in Appendix C.3.
@@ 7.2 Global n-gram rescoring
Figure 7: Block efficiency of $p$-best n-gram BPD methods as a function of the number of top $p$ sequences verified in parallel. The block efficiency of the methods is evaluated with the the same number of paths extracted from the top-$16$ lattice.
@@ 7.3 Empirical evaluation
Block efficiencyTable 4 and Figure 7 demonstrate the impact of lattice rescoring on block efficiency across various tasks. Autoregressive neural, n-gram LM, and $p$-best n-gram BPD rescoring all demonstrate improvements in block efficiency, although gains are task-dependent.
  • High initial block efficiency (LAMBADA, CNN/Daily): Both rescoring methods show little to no improvement, suggesting that vanilla BPD already produces high quality drafts.
  • Low initial block efficiency (SQuAD V1, SAMSUM, XSUM, NewsRoom): Both neural and n-gram rescoring lead to block efficiency gains, particularly with neural LMs achieving the best performance in some cases. This suggests that rescoring helps refine predictions and navigate the lattice more efficiently in these scenarios.
Table 5: Wins, ties, and losses of 61M parameter neural-rescored drafts and vanilla drafts. “Repair” corresponds to instances where the rescored draft eliminates repetition and “Regress” corresponds to instances where the rescored draft introduces repetition.
@@
Repairing repetitions  In subsection 6.1, we note that vanilla block drafts are prone to token-level repetition and that rescoring with a simple language model reduces the incidence of this. Although rescoring reduces repetition overall in drafts, is this driving improvements in block efficiency? To answer this, we compared the drafts generated by greedy rescoring with the 61M parameter neural rescorer against vanilla drafts. Time step instances were considered wins/ties/losses based on the accepted prefix length of the rescored draft vs. vanilla draft. Table 5 displays the win frequency across tasks along with the percentage of wins/losses attributed to introducing/eliminating repetition.
Note that in the tasks where rescoring improves block efficiency the most, NewsRoom and MultiNews, a high percentage of those repaired instances are driven by fixing erroneously repeated tokens. In fact, for MultiNews, 66.23% of block drafts are improved through repetition repair. We also evaluated the performance of rescoring with in-domain trained rescoring LMs, but found that they tended to perform no better than C4-trained LMs (Appendix D).
@@ 8 Conclusion
This paper presents a comprehensive analysis of BPD, highlighting its predictive dynamics and proposing methods to refine the generation of block drafts. Our study offers insights into BPD’s behavior, particularly the tendency for drafts to contain consecutive repetitions and its heads to exhibit varying confidence levels in predictions. We introduce a novel measure, oracle top-$k$ block efficiency, to explore potential improvements in block efficiency. Two algorithms are proposed for generating higher quality drafts: one for local rescoring with small neural models (i.e., neural BPD) and another for global rescoring with an n-gram LM and generating multiple drafts (i.e. $p$-best n-gram BPD). These algorithms leverage the strengths of both blockwise parallel LMs and small rescoring models to reduce average decoding latency, pushing the boundaries of efficient text generation with BPD. We believe that this paper lays the groundwork for future exploration in optimizing LM decoding speed.
@@ Appendix A Broader Impact
  • Inference efficiency: We achieve further acceleration in inference with minimal increases in CPU overhead and memory bandwidth utilization. In particular, we show that our methods significantly boost the inference efficiency compared to the standard autoregressive inference.
  • Scalability: The introduction of selecting top $p$ sequences from top $k$ predictions demonstrates notable scalability, hinting at accelerated inference with growing model sizes. This scalability is further explored through diverse experiments involving various n-gram  and neural LM architectures.
  • Systematic impact: The advancement of BPD promises identical generation quality to traditional autoregressive methods, but with significantly improved efficiency in decoding times, which can be helpful in latency-critical applications.
@@ B.1 Limitation
In this work, we focus on refining block drafts for greedy decoding, the decoding scheme considered in the original BPD paper [38]. Extending this methodology to non-greedy sampling strategies is an area for future work. Leviathan et al. [24] demonstrates the feasibility of applying speculative decoding strategies to accelerate (non-greedy) sampling. Their approach underscores the universality of the verification process across diverse model architectures, contingent on the availability of draft token-level predictions and probabilities — a requirement met by (rescored) block candidates.
@@
We show that lattice rescoring can improve the block efficiency for a 1.5 billion parameter blockwise parallel LM. It remains an open question as to whether one would observe similar improvements in block efficiency with larger LMs.
@@
A notable challenge in our framework is the memory overhead encountered during parallel verification. The simultaneous processing of multiple token predictions, particularly for larger models and draft batches when coupled with lattice rescoring, can introduce high demands on memory. That said, this can be mitigated by the application of tree attention during verification [37].
@@
The implementation of drafting heads within our BPD framework, while hewing close to the original BPD proposal, leaves room for architecture experimentation. The choice of architecture for the block drafter is crucial, and will likely affect the gains one observes from rescoring the block draft top-$k$ lattice. Future iterations could benefit from more sophisticated training methods, enhancing these heads’ ability to navigate complex linguistic contexts and improving overall predictive performance.